OpenStudy (yrivers36):

Can someone help out with this problem please. x^2-3x=0 with explanation

OpenStudy (anonymous):

x^2=3x so x^2/x=3 So x=3

OpenStudy (yrivers36):

How did you get x=3

OpenStudy (anonymous):

since x^2-3x =0, x^2=3x

OpenStudy (anonymous):

so just arrange x to one side you will get x^2/x=3

OpenStudy (yrivers36):

divide x^2 by 3x

OpenStudy (anonymous):

or check your answer like this 3^2-(3)(3)=0

OpenStudy (yrivers36):

I know that the answer is x=3 because it comes out to 9-9=0 but I dont know how you got rid of the x^2 to get x=3

OpenStudy (anonymous):

x times x divided by x

OpenStudy (yrivers36):

Thank you I understand

OpenStudy (anonymous):

I can't type square in my system that's why used '^'